A man bought 9 pounds of coffee and 6 pounds of butter for a total of $47.75
A day later, the prices had not changed and he bought 10 pounds of coffee and 14 pounds of butter for $64.00
Find the price per pound of the butter.

Answer by mananth(16949)

coffee x $/kg
butter y $/kg

9 x + 6 y = 47.75 .............1
10 x + 14 y = 64 .............2
Eliminate y
multiply (1)by -14
Multiply (2) by 6
-126 x -84 y = -668.5
60 x 84 y = 384
Add the two equations
-66 x = -284.5
/ -66
x = 4.31 OR 4 23/74
plug value of x in (1)
9 x + 6 y = 47.75
38.8 + 6 y = 47.75
6 y = 47.75 -38.8
6 y = 8.95
y = 1.49
coffee 4 1/3 $/kg
butter 1 1/2 $/kg
m.ananth@hotmail.ca